In a group of 42 students, 22 take history, 17 take biology and 8 students took both history and biology. Find the probability of a randomly selected student

a) Took both biology and statistics?

b)Took biology, but not history?

(In question a. my paper says statistics, it is not part of the given.)
Answer #1

a) P(a student took both biology and statistics) = Number of students who took both / Total number of students

= 8/42

= 0.1905

b) Number of students who took biology but not history = 17 - 8 = 9

P(a student too biology, but not history) = 9/42

= 0.2143